OpenStudy (anonymous):

40 in.= ft what is the awnser?

OpenStudy (anonymous):

2 feet some inches

OpenStudy (anonymous):

\[\left\{\text{Floor}\left[\frac{40}{12}\right] \text{feet},\text{ }\text{Mod}[40,12] \text{ inches}\right\} = \{3 \text{ feet},4 \text{ inches}\}\] Floor is a function that returns only the quotient from a division. The Mod function returns the remainder from a division. These functions are provided with Mathematica and I assume they are built into some of the more advanced hand held calculators.